OpenStudy (micahm):

Solve each equation, if possible. Write irrational numbers in simplest radical form. Describe the strategy you used to get your solution and tell why you chose that strategy. a. x^2+4=0 b.x^2-6x+1=0 can i have step by step help

OpenStudy (solomonzelman):

you can already see that if you subtract 4 from both sides (in the first equation), you won't get any real solutions.

OpenStudy (micahm):

ok

OpenStudy (solomonzelman):

have you learned about imaginary number \(\large\color{black}{i }\) (which is equivalent to \(\large\color{black}{ \sqrt{-1} }\) ) ?

OpenStudy (micahm):

ok

OpenStudy (solomonzelman):

is that a yes?

OpenStudy (micahm):

my teacher told me this Comments: You haven’t solved either equation. You want to take each part separately. If you would redo part B only and resubmit for grading to me. You will want to use the quadratic formula.

OpenStudy (solomonzelman):

i would complete the square for equation 2.

OpenStudy (solomonzelman):

\(\large\color{black}{ x^2-6x+1=0}\) \(\large\color{black}{ x^2-6x+1\color{red}{+8}=0\color{red}{+8}}\) \(\large\color{black}{ x^2-6x+9=8}\) and on....

OpenStudy (micahm):

so A does not need to be solved at all

OpenStudy (micahm):

so the problem continue to the answer 8